You bought a stock three months ago for $34.18 per share. The stock paid no dividends. The current share price is $35.07.

What is the APR and EAR of your investment?

Please explain your answer in detail. Thanks!
Answer #1

APR is annual percentage rate

calculate 3 month rate increment

35.07 = 34.18*(1+r)

r = 2.6038% (for 3 month)

APR = 2.6038*4 = 10.4152%

EAR = 1.026038^4-1 = 10.8290% (it takes into a/c the compounding)